[Hallc_running] [Revised Logentry] RC daily update

simona at jlab.org simona at jlab.org
Thu Feb 1 11:45:02 EST 2018

Logentry Text:
Yesterday at 3:30 pm we finished the 5 pass commissioning running and took an access to insert the GEM in front of the SHMS HB. We ran with the GEM inserted for a while at ~0.3 muA then we took another access to retract it. I assume Latif will post a summary of the GEM data taking. Thanks to Jerry (Hall C tech) for supporting this configuration change.

We then started to set the spectrometers for the H(e,e'p) check. This involved changing polarity for the SHMS to detect protons and also switching the DAQ from single arm to coincidence mode. The SHMS magnets were very well behaved. The HMS ones not so much - it's not always possible to go to the cycling currents for the quads. 
The dipole is still set by current make adjustments until the NMR reads something very close to the desired field setting. Sometime during the swing shift the HMS dipole tripped: the power supply turned off and went into local. We switched it to remote from the counting house and ramped up the dipole again.

We took H(e,e'p) data until middle of the owl shift when we switched to physics production for the CT experiment. We took LH2 data in the morning and then switched to C 6%. We will be doing production on this target for a while.

Of note:

--> beam current as delivered so far to Hall C is not what the experiment needs: for most of the swing and owl shifts we were limited to 15 to 25 muA with very brief periods of 35 to 40 muA (but higher trip rate). Color Transparency needs at least 40 to 50 muA. 
My understanding of Jay's and others take on what might be going on: there is a beam current loss in the machine that's detectable in both A and C and there is a suspicion that it's real and not induced by the instrumentation that's measuring it (sensitivity of BCM cavities to temperature) - or at least not all of it.  Jay talks about an apparent 1 muA loss which is preventing high current beam delivery to Hall C. After our desperate cries for higher current, the accelerator delivered 45 muA throughout the morning but only by changing this limit on the beam current loss. 
For more on the beam current loss issue check: https://logbooks.jlab.org/entry/3525103

Beam to Hall C has been interrupted this morning for Jay's investigation at the Lambertson (https://logbooks.jlab.org/entry/3525186 and https://logbooks.jlab.org/entry/3525214). Beam delivery to Hall C was resumed at about 10:55 am.

--> this morning we were alerted by Mike F. that the HMS dipole LHe supply JT 5 valve (top fill valve) could not maintain an appropriate He level  (it dropped to something like 40 %): https://logbooks.jlab.org/entry/3525164. There was talk about going in the Hall to troubleshoot and fix the problem but the magnet group decided in the end to wait until Tuesday for the access (beam studies). Meanwhile the JT 1 valve (bottom fill valve) will be used to help maintain the LHe level for the dipole. We really need to keep an eye on the He levels. Check Thia's entry for more details: https://logbooks.jlab.org/entry/3525210

This morning I asked Steve Wood to add the cryo levels for the magnets into the alarm handler. Hopefully this will be done today. Regardless, we should keep the HMS dipole valves screen open at all times and monitor the status periodically as per Thia's entry (https://logbooks.jlab.org/entry/3525210)

We will meet today at 1:30 pm on the second floor of the counting house.

Link: https://bluejeans.com/543903672
Connecting directly from a room system?
1) Dial: or bjn.vc
2) Enter Meeting ID: 543 903 672
Just want to dial in on your phone?
1) Direct-dial with my iPhone or
+1.408.740.7256 (United States)
+1.888.240.2560 (US Toll Free)
+1.408.317.9253 (Alternate number)
(Global Numbers)
2) Enter Meeting ID: 543 903 672
3) Press #


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3525219
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180201/996cc55b/attachment.html>

More information about the Hallc_running mailing list